OpenStudy (henryarias14):

CALCULUS derivative of : f(x)=2Arcsin(x-1) please help me through the steps

OpenStudy (sunnnystrong):

Okay so: Recall that- \[\frac{ d }{ dx }\arcsin(x)=\frac{ 1 }{ \sqrt{1-x^2} }\frac{ d }{ dx }\]

OpenStudy (henryarias14):

yes

OpenStudy (sunnnystrong):

This is a chain rule --> So you are going to take the derivative of the outside function first (arcsin) leaving the inside alone and multiplying it by the derivative of the inside...

OpenStudy (henryarias14):

okay so \[\frac{ 1 }{ \sqrt{1-x^2} } *2 \] ?

OpenStudy (henryarias14):

the book gives me that answer and thats what i got but under the square root it gives me 2x-x^2 and idk how they got that 2x

OpenStudy (sunnnystrong):

@henryarias14 yep... but what is x? It's (x-1) right?

OpenStudy (henryarias14):

yes

OpenStudy (henryarias14):

wait. so sqr(1-(x-1)) ?

OpenStudy (sunnnystrong):

\[f'(x)=\frac{ 2 }{ \sqrt{1-(x-1)^2} }\]

OpenStudy (henryarias14):

ohhhh

OpenStudy (sunnnystrong):

yep:) and we can simplify it further

OpenStudy (henryarias14):

i get sqr(x^2+2x) but its still not what the book tells me. it says it should be 2x-x^2 unless i did something wrong in my notes. what did you get?

OpenStudy (sunnnystrong):

\[f'(x)=\frac{ 2 }{ \sqrt{1-(x^2-2x+1)} }\] *Expand and cancel -1 \[f'(x)=\frac{ 2 }{ \sqrt{2x-x^2} }\]

OpenStudy (henryarias14):

oh i get it now. thanks!!!

OpenStudy (henryarias14):

got a final on thursday

OpenStudy (sunnnystrong):

*Multiply by radical \[f'(x)=\frac{ 2 }{ \sqrt{2x-x^2} }\] \[f'(x)=\frac{ 2\sqrt{2x-x^2} }{ 2x-x^2 }\] & @henryarias14 :D good i'm happy!

OpenStudy (henryarias14):

oh yeah forgot about the radical. thanks! although i think my professor doesn't mind if we leave the radical on the denominator for this type of problems lol

OpenStudy (sunnnystrong):

Yeah they usually won't care lol :P